# Can LLM “Self-report”?: Evaluating the Validity of Self-report Scales in Measuring Personality Design in LLM-based Chatbots

### v1

@@ 2.4 Construct Validity Evaluation
  • Convergent validity assesses whether different methods measuring the same trait yield consistent results John and Benet-Martínez (2000). We tested this by comparing the chatbot’s self-reported personality scores across three established scales. Strong correlations among these measures indicate consistent measurement across different methods.
@@
  • Discriminant validity evaluates the distinctiveness of different traits Campbell and Fiske (1959). We verified the independence of the chatbot’s personality dimensions by calculating correlations between different trait scores. Low correlations between unrelated traits confirm good discriminant validity.
@@
  • Criterion validity examines the correlation between a measure and an external criterion Cronbach and Meehl (1955). We assessed this by comparing the chatbot’s self-reported personality traits with human perceptions, measured through participant ratings after interacting with the chatbot. The value of criterion validity indicates the level of alignment between self-reports and user perceptions.
@@
  • Predictive validity refers to a measure’s ability to predict future outcomes or behaviors Funder (2006). We evaluated this by exploring how the chatbot’s self-report personality influences the quality of interaction, as captured through the UEQ. A strong relationship between personality traits and positive user experiences suggests that these traits are valid predictors of chatbot effectiveness.
@@
To assess convergent and discriminant validity, we applied the Multitrait-Multimethod (MTMM) matrix approach Campbell and Fiske (1959). This framework evaluates multiple traits across different methods to identify patterns of consistency and distinctiveness. We implemented this analysis using Spearman correlation in R. Due to the non-hierarchical structure of the BFI-2-XS, which does not decompose traits into facets, we focused on domain-level traits to evaluate construct validity across different assessment settings.
@@ 3 Results
3.1 Does Personality Setting Work?
Domain Job Interview Public Service Social Support Travel Planning Guided Learning
@@ 3.1 Does Personality Setting Work?
Ope 3.43 0.57 3.23 0.83 3.80 0.47 3.13 0.89 3.83 0.50 3.23 0.94 4.00 1.04 3.27 0.70 3.70 1.07 3.57 0.61
Table 1: Domain-level mean (M) and standard deviation (SD) of human-perceived personality scores under high and low personality settings across five tasks. Light gray cells indicate instances where the low standard deviation for high or low personality settings and the difference between these settings exceeds 0.3. Dark gray cells highlight instances where the same difference exceeds 0.5.
Note: n = 500 in total. Ratings were made on a five-point scale ranging from 1 (disagree strongly) to 5 (agree strongly). Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
We compared the mean and standard deviation of task-specific human-perceived personality evaluation scores across different levels to assess how effective our personality design method is. As shown in Table 1, with the exception of conscientiousness scores in the social support task, all other domains consistently score higher in the high-setting conditions across tasks, with moderate variances. Additionally, the analysis of variance (Appendix G, Table 10) indicates that the variance between high-low groups is more pronounced than the variance within each group. These findings collectively support the effectiveness of our trait manipulation in the chatbot personality design.
3.2 Convergent and Discriminant Validity
Table 2 presents the correlation between the chatbot’s self-reported scores across different scales. It is evident that, regardless of the dimension, the correlations across scales show a high degree of consistency, with an average correlation coefficient of 0.85. This result indicates that the chatbot demonstrates a high level of stability in its self-reports across different personality questionnaires, suggesting consistency within self-report methods.
@@ 3.2 Convergent and Discriminant Validity
BFI-2-XS vs IPIP 0.82 0.84 0.91 0.86 0.81 0.85
IPIP vs BFI-2 0.88 0.83 0.86 0.86 0.83 0.85
Table 2: Correlation between self-reported personality scores using BFI-2-XS, BFI-2, and IPIP-NEO-120.
Note: n = 500 in total. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
Table 3: Correlation analysis for human-perceived and self-report personality scores across personality traits.
Note: n = 500 in total. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
Table 3 shows the correlations of human-perceived and self-report personality scores across different personality traits. The results further demonstrate alignment across the self-report methods, both in terms of the magnitude and direction of correlations between each pair of traits, as well as the absolute mean values.
3.3 Criterion Validity (Self-report vs. Human)
A good measure for personality design should align with the end-user perception of the chatbot. Table 4 presents the means, standard deviations, and correlations between human perception and chatbot self-reported scores. The results show that the means of the human perception scores and chatbot self-reported scores are quite similar, primarily reflecting the overall mean distribution of the sample, which is expected to be close. Notably, under the same personality setting, the standard deviation of the human perception scores is smaller than that of the chatbot self-reported score, indicating that human perceived scores show less variation between individuals, while chatbot self-reports exhibit greater variability.
What is more interesting is the correlation between human perception scores and chatbot self-reported scores. Apart from the relatively high correlation (0.58 ± 0.02) in the domain of agreeableness, the correlations in the others are all below 0.5. This suggests a low level of consistency between human perceptions and chatbot self-reports.
In addition, Table 3 shows differences in discriminant validity between human-perceived personality and chatbot’s self-reported personality. Compared to human-perceived personality traits, chatbot’s “self-reported” traits show a higher correlation with each other, except for the correlation between conscientiousness and openness, suggesting a lower level of discriminant validity.
Human-perceived BFI-2-XS BFI-2 IPIP-NEO-120 Self-reported Mean
@@
Con 3.97 0.96 3.70 1.42 0.23 3.44 1.10 0.24 3.62 1.12 0.25 3.59 1.21 0.24
Neu 2.03 0.88 2.30 1.30 0.29 2.71 1.11 0.33 2.37 1.12 0.36 2.46 1.18 0.33
Ope 3.27 0.84 3.31 1.32 0.23 3.09 1.11 0.21 3.36 0.83 0.24 3.25 1.09 0.23
Table 4: Domain-level mean (M), standard deviation (SD), and correlation ($\rho$) of self-reported and human-perceived personality scores using the BFI-2-XS, BFI-2, and IPIP-NEO-120.
Note: n = 500 in total. Ratings were made on a five-point scale ranging from 1 (disagree strongly) to 5 (agree strongly). Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
Across the Tasks
In addition to the overall correlation analysis, we further examined each individual task. Given the high correlations among the three “self-report” scales (§3.2), we averaged the results of the three questionnaires and conducted a comparative analysis with human-perceived personality design (see Table 5). For detailed correlation analysis for each questionnaire, please refer to Tables 11, 12 and 13 in Appendix H.
As shown in the table, agreeableness consistently exhibited the highest correlation across all tasks. However, the correlations for other personality dimensions were weaker, with average values generally below 0.39, and considerable fluctuations in the correlations for the same personality traits across different tasks. For instance, conscientiousness even showed a negative correlation in the social support task, further highlighting the limitations of evaluating chatbot personality design solely based on its “self-report” data, especially across different task contexts.
One key factor may explain these results. The variation in correlation across tasks for specific personality traits may reflect how the chatbot’s behavior is influenced by the task type and the interaction with users. For example, agreeableness may align more with general expectations of friendliness and cooperation during interactions, whereas conscientiousness might have the opposite effect in some tasks (such as social support tasks) due to an excessive focus on detail or rules, which could negatively impact the human experience. The performance of different personality traits across tasks is not uniform but shaped by task context and interaction.
@@
Table 5: Average correlation between self-reported and human-perceived personality scores across tasks.
Note: n = 100 for each task. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
In summary, the criterion validity of LLM-based chatbot is low, raising concerns about the effectiveness of using “self-report” methods to measure a chatbot’s personality design. Such inconsistency indicates the gap between an LLM’s responses to personality questionnaires and the actual behaviors it exhibits during real-world user interactions.
3.4 Predictive Validity (Interaction Quality)
Task Ext Agr Con Neu Ope
Job Interview 0.26 0.37 0.49 -0.40 0.48
Public Service 0.14 0.32 0.41 -0.22 0.28
Social Support -0.05 0.17 0.19 -0.15 0.25
Travel Planning 0.28 0.71 0.76 -0.55 0.52
Guided Learning 0.15 0.14 0.24 -0.09 0.16
@@ 3.4 Predictive Validity (Interaction Quality)
Table 6: Correlation between aggregated UEQ scores and human-perceived personality scores across tasks.
Note: n = 433 in total. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
Tables 6 and 7 detail correlation analyses between aggregated UEQ scores and personality scores, both human-perceived and self-reported, across five tasks. This analysis evaluates the predictive validity of these personality assessments in relation to interaction quality with chatbots, which is the ultimate goal of building such chatbots. Table 6 demonstrates significant correlations between perceived agreeableness and conscientiousness and user experience, notably in the travel planning task, with correlations of 0.71 and 0.76, respectively. These findings suggest that chatbots perceived as agreeable and conscientious markedly enhance user experience. In contrast, neurotic traits show consistently negative correlations across all tasks, most pronounced with a correlation of -0.55 in travel planning, indicating that these traits detract from usability. The data illustrates that positive personality traits are predictive of enhanced usability, while negative traits hinder it. This consists with research indicating the crucial role of personality design in shaping user experience among the selected tasks.
Task Ext Agr Con Neu Ope
Job Interview 0.12 0.15 0.13 -0.09 0.03
Public Service 0.03 0.04 0.17 -0.16 -0.03
Social Support 0.10 0.10 0.08 -0.10 0.02
Travel Planning 0.03 0.08 0.04 -0.17 -0.09
Guided Learning -0.06 -0.04 -0.07 0.02 -0.01
Table 7: Correlation between aggregated UEQ scores and self-reported personality scores across tasks.
Note: n = 433. Ext = Extraversion; Agr = Agreeableness; Con = Conscientiousness; Neu = Neuroticism; Ope = Openness.
Conversely, Table 7 reveals discrepancies between self-reported personality scores and user experience, characterized by low and inconsistent correlations. For instance, conscientiousness exhibits a modest positive correlation of 0.17 in the public service task, the highest observed among the traits. However, the overall pattern of weak and occasionally near zero correlations, such as -0.01 for openness in guided learning, indicates that self-reported traits are unreliable predictors of user experience. This disparity likely arises because self-report measures fail to capture the dynamic impacts of interaction and task setting, crucial in shaping user perceptions during real-time chatbot engagements. These findings emphasize the limitations of exclusively relying on self-reported personality to evaluate user experience and the importance of considering perceived personality in interactive settings. Future research should explore the factors affecting the relationship between personality perception and user experience, and develop methods that more accurately capture the dynamic nature of personality in interactions.
In summary, self-report personality scales failed to correlate with interaction quality, which indicates a disconnect between the model’s response to personality items and how their behaviors manifest during real interaction. It highlights the complexity of user interaction and the challenges in evaluating personality design in LLM-based chatbots. As the ultimate goal of personality design is to improve interaction, such a validity link is crucially important. The validity issues of the “self-report” evaluation method may misguide chatbot development.
4 Towards Interaction and Task Grounded Personality Evaluation
@@
Current self-report scales assume that personality traits are expressed consistently across different scenarios. However, the limited predictive and criterion validity of self-report personality scales, as shown in our findings, suggests a disconnect between the scores and the user experience. Moving forward, we advocate for transitioning from static, questionnaire-based evaluations to task-driven assessments that better reflect the scenarios where chatbots operate, aligning with calls from prior research Lee et al. (2022); Liao and Xiao (2023).
@@
First, personality evaluations should be based on specific tasks or scenarios, as chatbot personality traits manifest differently depending on the situation, similar to how humans adjust their behavior based on context Sauerberger and Funder (2017). Furthermore, contextual shifts may lead users to form different assumptions about a chatbot’s personality, as social roles influence how individuals attribute traits to others Roberts (2007). Thus, evaluation methods should account for biases from user expectations and the chatbot’s context-driven behaviors to align with design objectives.
@@
Second, personality evaluations without considering the expression of traits in real-world interactions fail to capture the chatbot’s impact on user experience. Since personality is conveyed through behaviors and mutual perceptions in interaction Geukes et al. (2019), an ideal evaluation framework should account for factors in continuous interactions, such as response patterns, and adaptability to user input. To understand the patterns essential for user experience, we may observe how user perceptions of personality evolve with interactions and how these changes impact user satisfaction. By embracing task-based interactive evaluation, we can improve personality evaluations, ultimately fostering chatbots that enhance user satisfaction.
5 Related Work
Human-Like Responses in Large Language Models
Recent studies have shown that LLMs demonstrate remarkable abilities resembling human-like responses. For example, Lampinen et al. (2024) found that LLMs display response patterns similar to humans in logical reasoning tasks, with similar findings being reported by Kosinski (2023) and Wang et al. (2024a). Additionally, Pellert et al. (2023) found that LLMs perform comparably to humans across various psychological scales, such as value orientations and moral norms. All these abilities have garnered significant interest from both computer scientists and social scientists, as these findings suggest that we may be able to draw upon existing social science research on humans as a reference for LLM studies. For instance, personality, which has been extensively studied in psychology and consistently proven to be a reliable predictor of various behavioral and psychological outcomes, is one such area of interest. Personality encodes rich and complex information in textual data  Goldberg (1990); Saucier and Goldberg (2001), so LLMs may capture and model such encoding by learning from vast training data. Many LLM studies explore the personality traits these models exhibit, with the hope that we can similarly use the models’ simulation of personality to predict or shape their behavior Lee et al. (2024); Huang et al. (2023b); Wang et al. (2024b); Ran et al. (2024).
Current Approaches for Language Models Personality Assessment
Currently, most measurements of LLMs’ personality are based on psychological research related to human personality assessment, typically using personality scales. A common approach is to have the model respond to scales that have already been developed and validated on human samples, such as the BFI-2 Soto and John (2017a) and HEXACO Lee and Ashton (2004, 2006). The model’s scores are then calculated based on its answers to the multiple-choice questions (MCQs) on these scales. The rationale behind this is straightforward: since we believe that LMs simulate human abilities, we can directly borrow well-established and validated methods used for humans to assess the personality traits of these models. For example, studies like Huang et al. (2023b) used this approach to evaluate whether LLMs reflect the personality traits outlined in the BFI John et al. (1999). Serapio-García et al. (2023) presented the reliability and validity of LLMs in resembling personality profiles along desired dimensions with IPIP-NEO Goldberg et al. (1999) and BFI John et al. (1999). However, some researchers argue that traditional scales may not sufficiently capture the personality of LMs and have adopted alternative approaches. For instance, some researchers modify scale questions into prompts that elicit text-based responses Wang et al. (2024c), or others have experts rate stories generated by LMs according to specified criteria Jiang et al. (2023). These text-based evaluation methods may better reflect real-world application scenarios.
Existing methods have certain shortcomings, as they often overlook the importance of analyzing model behavior from the perspective of the task-specific performance. In other words, current approaches may overemphasize on directly measuring the model’s predefined personality traits, without adequately considering the model’s specific performance in actual tasks. If the ultimate goal of defining a model’s personality is to enhance its effectiveness in specific tasks, then the behavior exhibited by the model in these tasks—and the corresponding personality traits—should be the primary focus of research and evaluation.

### v2

@@ 3.4 Analysis Method
  • Convergent validity assesses whether different methods measuring the same trait yield consistent results (John & Benet-Mart´ınez, 2000). To assess this, we compared the chatbot’s self-report personality scores across three established scales. Strong inter-scale correlations indicate consistent responses under our personality settings.
@@
  • Discriminant validity evaluates the distinctiveness of different traits (Campbell & Fiske, 1959). We calculated correlations between different personality trait scores across psychometric scales and evaluation methods. Low correlations between unrelated traits confirm good discriminant validity.
@@
  • Criterion validity examines the correlation between a measure and an external criterion (Cronbach & Meehl, 1955). We assessed this by comparing the chatbot’s self-report personality traits with human perceptions, measured through participant ratings after interacting with the chatbot. The value of criterion validity indicates the level of alignment between self-reports and human perceptions.
@@
  • Predictive validity refers to a measure’s ability to predict future outcomes or behaviors (Funder, 2006). We evaluated this by examining whether the chatbot personality setting predicted the quality of interaction, as captured through the UEQ. Strong correlations between personality evaluation scores and positive user experiences suggest that personality traits are valid predictors of chatbot effectiveness.
@@
To assess convergent and discriminant validity, we employed the Multitrait-Multimethod (MTMM) matrix approach (Campbell & Fiske, 1959), which evaluates multiple traits across different methods to identify patterns of consistency and distinctiveness. We conducted this analysis using Spearman correlations. Given the non-hierarchical structure of the BFI-2-XS, our analysis focused on domain-level personality traits.
@@ 4 Results
4.1 Personality Settings Work for Both Human and Chatbot
Evaluation Domain Job Interview Public Service Social Support Travel Planning Guided Learning
@@ 4.1 Personality Settings Work for Both Human and Chatbot
OPE 3.43 0.57 3.23 0.83 3.80 0.48 3.07 0.95 3.83 0.50 3.23 0.94 4.00 1.04 3.27 0.70 3.70 1.07 3.57 0.61
Table 1: Domain-level mean (M) and standard deviation (SD) of self-report and human-perceived personality scores under high and low settings.
Personality setting work but chatbot’s self-report scores vary more significantly across personality conditions. To evaluate the effectiveness of our personality settings, we examined descriptive statistics of self-report and human-perceived scores. Specifically, we tested for significant between-group differences under high and low personality conditions, which indicates successful trait manipulation. As shown in Table 1, self-report scores are consistently higher in high-setting conditions across tasks. Human-perceived scores followed a similar trend, except for Conscientiousness in the social support task. Traits that are more readily conveyed through language and interaction style (Mairesse & Walker, 2010), such as Agreeableness, showed larger gaps between high and low personality conditions in human evaluations. Overall, the difference between conditions was more pronounced in self-reports than in human perceptions.
To further support this, we ran the analysis of variance to test whether the mean difference between the high and low conditions is statistically significant. This method compares between-group variance, which reflects differences across conditions, with within-group variance, which captures individual variability within each condition, to determine whether the means under different conditions diverge meaningfully. A significantly larger between-group variance indicates that the manipulation had a meaningful effect on the dependent variable, supporting the hypothesis that the group means differ. As shown in Table 9 in Appendix G, F values for the chatbot’s self-report scores exceed the conventional threshold of 1. Human-perceived scores show a similar pattern, with many F values also exceeding the threshold of 1; however, the magnitude of this effect is smaller relative to that observed for the self-report scores. We provide a more in-depth discussion of this difference in §4.2. Overall, these results collectively support the effectiveness of our personality manipulations.
Personality setting effects are consistent across psychometric scales. We further evaluated the chatbot’s self-report score on three personality inventories. Although these instruments differ in item content, they are designed to measure the same underlying traits. Consistent convergent and discriminant validity across them would indicate stable personality settings. Summarized in Table 2, the convergent validity results show that the chatbot’s self-report scores are highly consistent across the various inventories, with an average inter-scale correlation of over or equal to 0.85. Similarly, the discriminant validity analysis in Table 3 shows that correlations between distinct personality dimensions remain relatively uniform across scales, with a mean absolute correlation of approximately 0.53.
@@
BFI-2-XS vs IPIP-NEO-120 0.82 0.84 0.91 0.86 0.81 0.85
IPIP-NEO-120 vs BFI-2 0.88 0.83 0.86 0.86 0.83 0.85
Table 2: Correlation between self-report personality scores using different personality evaluation scales.
Trait Pair Human BFI-2-XS BFI-2 IPIP-NEO-120
[…]
Absolute Mean 0.39 0.54 0.54 0.51
Table 3: Correlation analysis for human-perceived and self-report personality scores across different personality traits.
4.2 Chatbot’s Self-report Scores Correlate with Human Perceptions Poorly
Personality measures differ in variability and validity. We believe that an effective measure for the chatbot’s personality design should align with the human perceptions of the chatbot. Table 4 presents the means, standard deviations, and correlations between human perception and chatbot self-report scores. As shown, the means across both sources are quite similar, suggesting general agreement at the aggregate level. However, most standard deviations of the human perception scores are smaller than that of the self-report scores, indicating that human ratings are more clustered while the chatbot’s self-reports show greater variability.
Human BFI-2-XS BFI-2 IPIP-NEO-120 Self-report Mean
@@ 4.2 Chatbot’s Self-report Scores Correlate with Human Perceptions Poorly
CON 3.98 0.96 3.70 1.42 0.24 3.44 1.10 0.25 3.62 1.12 0.25 3.59 1.21 0.25
NEU 2.03 0.88 2.30 1.30 0.29 2.71 1.11 0.33 2.37 1.12 0.36 2.46 1.17 0.33
OPE 3.27 0.83 3.31 1.32 0.24 3.09 1.11 0.21 3.36 0.83 0.24 3.25 1.09 0.23
Table 4: Domain-level mean (M), standard deviation (SD), and correlation ($\rho$) of self-report and human-perceived personality scores using different personality evaluation scales.
Statistical tests in Table 10 help explain this pattern. Compared to self-reports, human-perceived scores show smaller differences between high and low personality conditions, and greater variability across tasks. To further explore the differences between these two measures, we conducted paired t-tests within the same condition (Table 11). The significant p-values across all three psychometric scales reveal discrepancies between the two measures, suggesting that while self-report results may reflect intended personality settings, they do not align with how those traits are perceived during interactions. To control for potential false positives arising from multiple comparisons, we applied significance corrections (Table 12). Our conclusions remain unchanged after applying these adjustments, with most p-values remaining well below thresholds, confirming the robustness of the observed effects.
To further probe this discrepancy, we examine correlations between human perception and self-report scores. Apart from the relatively high mean correlation in the Agreeableness domain (0.58), the correlations in others are all below 0.4. This indicates a low level of consistency between how the chatbot evaluates its own personality and how humans perceive it, calling into question the reliability of self-report scores as indicators of perceived behavior. In addition, Table 3 shows differences in discriminant validity between the two measures. Compared to human perceptions, the self-report scores show a higher correlation across different personality traits, with the exception of the correlation between Conscientiousness and Openness. That suggests a lower level of discriminant validity among self-report methods than human perceptions.
Task context shapes personality expressions. To better understand how personality expression varies across different contexts, we analyzed the effect of task setting. Given the high correlations among the three self-report scales (§4.1), we averaged their results and compared them with human perceptions across individual tasks. As shown in Table 5, Agreeableness consistently had the highest correlation across tasks, while other traits exhibited weaker correlations and significant fluctuations across different tasks. For instance, in the social support task, Extraversion demonstrated near-zero correlation and Conscientiousness even showed a negative correlation, further highlighting that self-report measures are insufficient for capturing how personality traits emerge in the practical, interaction-based setting. A detailed breakdown by scales is provided in Table 14, Appendix I.
@@
Table 5: Average correlation between self-report and human-perceived personality scores.
A likely explanation for these results lies in the influence of task context on trait expression. Correlation variations across tasks may reflect how a chatbot’s behavior is influenced by the task type and the interaction with human. For example, Agreeableness may align more with general expectations of friendliness and cooperation during interactions. In contrast, Conscientiousness might have the opposite effect in some tasks, such as social support, where excessive focus on details or rules could negatively impact engagement. Personality traits’ performance across tasks is not uniform but shaped by task context and interaction.
In summary, these findings point to a disconnect between static self-report methods and dynamic human perception. The inconsistency between chatbot questionnaire responses and their observable behaviors in interactive contexts raises concerns about the reliability of using self-report measures alone to evaluate personality design in LLM-based chatbots.
4.3 Chatbot’s Self-report Scores Unreliably Predict User Experience
Table 6 summarizes the correlation between overall user-experience ratings and both human-perceived and self-report personality scores across the five tasks. This analysis assesses the predictive validity of the two personality-assessment approaches with respect to interaction quality. As indicated in Table 6, perceived Agreeableness and Conscientiousness exhibit robust positive relations with user experience, with a correlation of 0.71 and 0.76, respectively, in the travel-planning scenario. These magnitudes suggest that chatbots viewed as highly agreeable and conscientious substantially elevate perceived interaction quality. Conversely, Neuroticism demonstrates stable negative relations across all tasks, with the strongest negative correlation (-0.55) emerging in travel planning, implying that neurotic traits systematically undermine usability. Collectively, these correlation patterns underscore that desirable personality attributes predict superior usability, whereas undesirable attributes impede it, supporting prior work on the pivotal role of personality cues in shaping user experiences with conversational interfaces.
Task Evaluation EXT AGR CON NEU OPE
[…]
@@ 4.3 Chatbot’s Self-report Scores Unreliably Predict User Experience
Table 6: Correlations between UEQ scores and personality scores from self-report and human-perceived evaluations.
Conversely, the chatbot’s self-report data reveal only weak and inconsistent associations between trait scores and user experience. The strongest observed correlation, Agreeableness in the social support task, reaches only a modest correlation of 0.19. Across tasks, many relations are negligible or null (e.g., Openness and user satisfaction in the guided learning task, $\rho$ =-0.01), indicating that self-report traits are generally poor predictors of interaction quality. A plausible explanation is that standard self-report inventories fail to capture the situational and dialogic cues that shape real-time perceptions of a conversational agent. Accordingly, the present findings caution against relying solely on dispositional self-reports to gauge user experience and highlight the added value of measuring perceived (interaction-level) personality. Future research should identify contextual moderators of the personality–experience linkage and devise assessment approaches that reflect personality as an emergent, dynamically co-constructed phenomenon within human-chatbot exchanges.
In summary, the absence of a strong correlation between self-report personality scores and objectively rated interaction quality points to a substantive disjunction between the traits reflected in questionnaires and the chatbot’s actual conversational behavior. This divergence highlights the multifaceted nature of user–agent exchanges and underscores the methodological challenges in assessing personality design for LLM-based chatbots. Since the ultimate purpose of endowing a chatbot with personality is to enhance user experience, establishing robust predictive validity between personality metrics and interaction outcomes is indispensable. Relying solely on self-report instruments risks misleading conclusions and may hinder effective system development.
5 Towards Interactive and Task Grounded Personality Evaluation
@@
Current self-report scales assume personality traits are expressed consistently across scenarios. However, our findings show that self-report scores exhibit limited predictive and criterion validity, suggesting a disconnect between the scores and the user experience.
To further investigate the role of interaction dynamics in developing personality evaluation methods that reflect real-world interaction scenarios, we fine-tuned GPT-4o (OpenAI, 2024) using all collected human-chatbot conversational transcripts and their corresponding human-perceived personality scores. As shown in Table 16, the personality scores generated by the fine-tuned model exhibit stronger correlations with human-perceived scores than self-report methods on average. This result suggests that integrating multi-turn contextualized conversation into the evaluation process can provide a more behaviorally grounded assessment, particularly in contexts where the alignment between system behavior and user impression is paramount. Detailed experimental settings are provided in Appendix J.
Building on these insights, we advocate for transitioning from static, questionnaire-based evaluations to task-driven assessments that better reflect the scenarios where chatbots operate, aligning with calls from prior research (Lee et al., 2022; Liao & Xiao, 2023).
@@
First, personality evaluations should be based on specific tasks or scenarios, as chatbot personality traits manifest differently depending on the situation, similar to how humans adjust their behavior based on context (Sauerberger & Funder, 2017). Furthermore, contextual shifts may lead users to form different assumptions about a chatbot’s personality, as social roles influence how individuals attribute traits to others (Roberts, 2007). Thus, evaluation methods should account for biases from in-context user expectations and the chatbot’s context-driven behaviors to align with design objectives.
@@
Second, personality evaluations that neglect the expression of traits in real-world interactions fail to capture the chatbot’s impact on user experience. Since personality is conveyed through behaviors and perceptions in interaction (Geukes et al., 2019), evaluation should consider factors in continuous interactions, such as response patterns and adaptability to input. This aligns with the Realistic Accuracy Model (RAM) (Funder, 1995), which emphasizes inferring personality traits from observable behavioral cues. To understand the patterns essential for user experience, we may observe how human perceptions evolve with interactions and how these changes impact satisfaction. By embracing task-based interactive evaluation reflecting human perceptions, we can improve chatbot design and enhance user satisfaction.